### v2

@@ 4 Declaration of competing interest
No. Corresponding quotient matrix $Q^{\pi}$
Characteristic polynomial $P_{Q}(G,x)$
$G_{3}$ $\left(\begin{array}[]{cccc}1&1&0&0\\ t&t+1&1&0\\ 0&1&3&2\\ 0&0&1&3\end{array}\right)$
$[x^{4}-(t+8)x^{3}+(6t+19)x^{2}-(7t+16)x+4](x-1)^{t}$
$G_{15}$ $\left(\begin{array}[]{ccccc}1&1&0&0&0\\ t&t+3&2&0&1\\ 0&1&3&0&0\\ 0&0&0&1&1\\ 0&1&0&1&2\end{array}\right)$
$[x^{5}-(t+10)+x^{4}+(6t+34)x^{3}-(10t+48)x^{2}+(3t+27)x-4](x-1)^{t}$
$G_{16}$ $\left(\begin{array}[]{cccc}1&1&0&0\\ t&t+4&2&2\\ 0&1&3&0\\ 0&1&0&3\end{array}\right)$
$[x^{3}-(t+8)x^{2}+(3t+15)x-8](x-3)(x-1)^{t+1}$
$G_{18}$ $\left(\begin{array}[]{cccccc}1&1&0&0&0&0\\ t&t+2&1&0&1&0\\ 0&1&2&1&0&0\\ 0&0&1&2&1&0\\ 0&1&0&1&3&1\\ 0&0&0&0&1&1\end{array}\right)$
$[x^{4}-(t+10)x^{3}+(7t+34)x^{2}-(13t+46)x+4t+20](x-1)^{t}x$
$G_{19}$ $\left(\begin{array}[]{cccccc}1&1&0&0&0&0\\ t&t+2&1&0&0&1\\ 0&1&2&1&0&0\\ 0&0&1&2&1&0\\ 0&0&0&1&2&1\\ 0&1&0&0&1&2\end{array}\right)$
$[x^{4}-(t+8)x^{3}+(5t+20)x^{2}-(5t+17)x+4](x^{2}-3x+1)(x-1)^{t-1}$
$G_{20}$ $\left(\begin{array}[]{cccccc}1&1&0&0&0&0\\ t&t+2&0&1&1&0\\ 0&0&1&1&0&0\\ 0&1&1&3&1&0\\ 0&1&0&1&3&1\\ 0&0&0&0&1&1\end{array}\right)$
$[x^{4}-(t+8)x^{3}+(5t+18)x^{2}-(3t+15)x+4](x^{2}-3x+1)(x-1)^{t-1}$
$G_{21}$ $\left(\begin{array}[]{cccccc}1&1&0&0&0&0\\ t&t+3&1&1&1&0\\ 0&1&2&1&0&0\\ 0&1&1&3&1&0\\ 0&1&0&1&3&1\\ 0&0&0&0&1&1\end{array}\right)$
$[x^{6}-(t+13)x^{5}+(9t+62)x^{4}-(26t+140)x^{3}+(27t+158)x^{2}-(8t+84)x+16](x-1)^{t-1}$
$G_{22}$ $\left(\begin{array}[]{cccccc}1&1&0&0&0&0\\ t&t+3&1&1&0&1\\ 0&1&2&1&0&0\\ 0&1&1&3&1&0\\ 0&0&0&1&2&1\\ 0&1&0&0&1&2\end{array}\right)$
$[x^{6}-(t+13)x^{5}+(9t+63)x^{4}-(27t+145)x^{3}+(31t+165)x^{2}-(11t+87)x+16](x-1)^{t-1}$
$G_{23}$ $\left(\begin{array}[]{cccccc}1&1&0&0&0&0\\ t&t+4&1&1&1&1\\ 0&1&2&1&0&0\\ 0&1&1&3&1&0\\ 0&1&0&1&3&1\\ 0&1&0&0&1&2\end{array}\right)$
$[x^{4}-(t+11)x^{3}+(6t+37)x^{2}-(7t+47)x+20](x-3)(x-1)^{t}$
$G_{26}$ $\left(\begin{array}[]{cccccc}1&1&0&0&0&0\\ t&t+2&1&0&0&1\\ 0&1&2&1&0&0\\ 0&0&1&3&1&1\\ 0&0&0&1&2&1\\ 0&1&0&1&1&3\end{array}\right)$
$[x^{5}-(t+12)x^{4}+(9t+51)x^{3}-(24t+94)x^{2}+(19t+71)x-16](x-1)^{t}$
$G_{27}$ $\left(\begin{array}[]{cccccc}1&1&0&0&0&0\\ t&t+2&0&1&0&1\\ 0&0&1&1&0&0\\ 0&1&1&4&1&1\\ 0&0&0&1&2&1\\ 0&1&0&1&1&3\end{array}\right)$
$[x^{6}-(t+13)x^{5}+(10t+61)x^{4}-(31t+135)x^{3}+(35t+150)x^{2}-(12t+80)x+16](x-1)^{t-1}$
$G_{28}$ $\left(\begin{array}[]{cccccc}1&1&0&0&0&0\\ t&t+3&1&1&0&1\\ 0&1&2&1&0&0\\ 0&1&1&4&1&1\\ 0&0&0&1&2&1\\ 0&1&0&1&1&3\end{array}\right)$
$[x^{6}-(t+15)x^{5}+(11t+84)x^{4}-(40t+228)x^{3}+(58t+319)x^{2}-(29t+221)x+60](x-1)^{t-1}$
$G_{29}$ $\left(\begin{array}[]{ccccc}1&1&0&0&0\\ t&t+3&1&0&2\\ 0&1&2&1&0\\ 0&0&1&3&2\\ 0&1&0&1&4\end{array}\right)$
$[x^{5}-(t+13)x^{4}+(9t+59)x^{3}-(23t+115)x^{2}+(16t+92)x-24](x-2)(x-1)^{t-1}$
$G_{30}$ $\left(\begin{array}[]{ccccc}1&1&0&0&0\\ t&t+3&0&1&2\\ 0&0&1&1&0\\ 0&1&1&4&2\\ 0&1&0&1&4\end{array}\right)$
$[x^{5}-(t+13)x^{4}+(9t+57)x^{3}-(21t+107)x^{2}+(10t+86)x-24](x-2)(x-1)^{t-1}$
$G_{31}$ $\left(\begin{array}[]{ccccc}1&1&0&0&0\\ t&t+4&1&1&2\\ 0&1&2&1&0\\ 0&1&1&4&2\\ 0&1&0&1&4\end{array}\right)$
$[x^{4}-(t+12)x^{3}+(7t+43)x^{2}-(8t+56)x+24](x-3)(x-2)(x-1)^{t-1}$
$G_{32}$ $\left(\begin{array}[]{cccccc}1&1&0&0&0&0\\ t&t+2&1&1&0&0\\ 0&1&3&1&0&1\\ 0&1&1&3&1&0\\ 0&0&0&1&2&1\\ 0&0&1&0&1&2\end{array}\right)$
$[x^{4}-(t+10)x^{3}+(7t+32)x^{2}-(11t+39)x+16](x^{2}-3x+1)(x-1)^{t-1}$
$G_{33}$ $\left(\begin{array}[]{cccc}1&1&0&0\\ t&t+2&0&2\\ 0&0&2&2\\ 0&1&2&3\end{array}\right)$
$[x^{3}-(t+8)x^{2}+(5t+17)x-2t-10](x-3)(x-2)(x-1)^{t-1}x$
$G_{34}$ $\left(\begin{array}[]{ccccc}1&1&0&0&0\\ t&t+3&1&2&0\\ 0&1&3&2&0\\ 0&1&1&3&1\\ 0&0&0&2&2\end{array}\right)$
$[x^{5}-(t+12)x^{4}+(8t+49)x^{3}-(17t+86)x^{2}+(8t+64)x-16](x-3)(x-1)^{t-1}$
$G_{35}$ $\left(\begin{array}[]{ccc}1&1&0\\ t&t+4&4\\ 0&1&5\end{array}\right)$
$[x^{3}-(t+10)x^{2}+(5t+25)x-16](x-3)^{2}(x-1)^{t}$
$G_{36}$ $\left(\begin{array}[]{cccccc}1&1&0&0&0&0\\ t&t+2&1&1&0&0\\ 0&1&4&1&1&1\\ 0&1&1&3&1&0\\ 0&0&1&1&3&1\\ 0&0&1&0&1&2\end{array}\right)$
$[x^{6}-(t+15)x^{5}+(12t+84)x^{4}-(48t+228)x^{3}+(77t+319)x^{2}-(41t+221)x+60](x-1)^{t-1}$
$G_{37}$ $\left(\begin{array}[]{cccc}1&1&0&0\\ t&t+2&0&2\\ 0&0&4&2\\ 0&1&2&3\end{array}\right)$
$[x^{4}-(t+10)x^{3}+(7t+29)x^{2}-(8t+28)x+8](x-3)(x-2)(x-1)^{t-1}$
$G_{38}$ $\left(\begin{array}[]{cccc}1&1&0&0\\ t&t+2&2&0\\ 0&1&5&2\\ 0&0&2&2\end{array}\right)$
$[x^{3}-(t+9)x^{2}+(6t+18)x-8](x-3)(x-2)(x-1)^{t}$
$G_{39}$ $\left(\begin{array}[]{ccccc}1&1&0&0&0\\ t&t+3&1&2&0\\ 0&1&4&2&1\\ 0&1&1&3&1\\ 0&0&1&2&3\end{array}\right)$
$[x^{5}-(t+14)x^{4}+(10t+68)x^{3}-(28t+146)x^{2}+(23t+139)x-48](x-3)(x-1)^{t-1}$
$G_{40}$ $\left(\begin{array}[]{ccccc}1&1&0&0&0\\ t&t+3&2&1&0\\ 0&1&5&1&1\\ 0&1&2&3&0\\ 0&0&2&0&2\end{array}\right)$
$[x^{5}-(t+14)x^{4}+(10t+67)x^{3}-(27t+142)x^{2}+(20t+136)x-48](x-3)(x-1)^{t-1}$
$G_{41}$ $\left(\begin{array}[]{cccc}1&1&0&0\\ t&t+4&2&2\\ 0&1&5&2\\ 0&1&2&3\end{array}\right)$
$[x^{4}-(t+13)x^{3}+(8t+51)x^{2}-(11t+75)x+36](x-3)^{2}(x-1)^{t-1}$
$G_{42}$ $\left(\begin{array}[]{cccc}1&1&0&0\\ t&t+2&2&0\\ 0&1&5&2\\ 0&0&2&4\end{array}\right)$
$[x^{4}-(t+12)x^{3}+(9t+43)x^{2}-(16t+56)x+24](x-3)(x-2)(x-1)^{t-1}$
$G_{43}$ $\left(\begin{array}[]{ccccc}1&1&0&0&0\\ t&t+3&2&1&0\\ 0&1&5&1&1\\ 0&1&2&4&1\\ 0&0&2&1&3\end{array}\right)$
$[x^{4}-(t+13)x^{3}+(9t+51)x^{2}-(15t+75)x+36](x-3)^{2}(x-1)^{t-1}$
$G_{44}$ $\left(\begin{array}[]{ccc}1&1&0\\ t&t+4&4\\ 0&1&7\end{array}\right)$
$[x^{3}-(t+12)x^{2}+(7t+35)x-24](x-3)^{3}(x-1)^{t-1}$

### v3

@@ 4 Declaration of competing interest
$G_{3}$ $\left(\begin{array}[]{cccc}1&1&0&0\\ t&t+1&1&0\\ 0&1&3&2\\ 0&0&1&3\end{array}\right)$ $[x^{4}-(t+8)x^{3}+(6t+19)x^{2}-(7t+16)x+4](x-1)^{t}$
$G_{15}$ $\left(\begin{array}[]{ccccc}1&1&0&0&0\\ t&t+3&2&0&1\\ 0&1&3&0&0\\ 0&0&0&1&1\\ 0&1&0&1&2\end{array}\right)$ $[x^{5}-(t+10)+x^{4}+(6t+34)x^{3}-(10t+48)x^{2}+(3t+27)x-4](x-1)^{t}$
$G_{16}$ $\left(\begin{array}[]{cccc}1&1&0&0\\ t&t+4&2&2\\ 0&1&3&0\\ 0&1&0&3\end{array}\right)$ $[x^{3}-(t+8)x^{2}+(3t+15)x-8](x-3)(x-1)^{t+1}$
$G_{18}$ $\left(\begin{array}[]{cccccc}1&1&0&0&0&0\\ t&t+2&1&0&1&0\\ 0&1&2&1&0&0\\ 0&0&1&2&1&0\\ 0&1&0&1&3&1\\ 0&0&0&0&1&1\end{array}\right)$ $[x^{4}-(t+10)x^{3}+(7t+34)x^{2}-(13t+46)x+4t+20](x-1)^{t}x$
$G_{19}$ $\left(\begin{array}[]{cccccc}1&1&0&0&0&0\\ t&t+2&1&0&0&1\\ 0&1&2&1&0&0\\ 0&0&1&2&1&0\\ 0&0&0&1&2&1\\ 0&1&0&0&1&2\end{array}\right)$ $[x^{4}-(t+8)x^{3}+(5t+20)x^{2}-(5t+17)x+4](x^{2}-3x+1)(x-1)^{t-1}$
$G_{20}$ $\left(\begin{array}[]{cccccc}1&1&0&0&0&0\\ t&t+2&0&1&1&0\\ 0&0&1&1&0&0\\ 0&1&1&3&1&0\\ 0&1&0&1&3&1\\ 0&0&0&0&1&1\end{array}\right)$ $[x^{4}-(t+8)x^{3}+(5t+18)x^{2}-(3t+15)x+4](x^{2}-3x+1)(x-1)^{t-1}$
$G_{21}$ $\left(\begin{array}[]{cccccc}1&1&0&0&0&0\\ t&t+3&1&1&1&0\\ 0&1&2&1&0&0\\ 0&1&1&3&1&0\\ 0&1&0&1&3&1\\ 0&0&0&0&1&1\end{array}\right)$ $[x^{6}-(t+13)x^{5}+(9t+62)x^{4}-(26t+140)x^{3}+(27t+158)x^{2}-(8t+84)x+16](x-1)^{t-1}$
$G_{22}$ $\left(\begin{array}[]{cccccc}1&1&0&0&0&0\\ t&t+3&1&1&0&1\\ 0&1&2&1&0&0\\ 0&1&1&3&1&0\\ 0&0&0&1&2&1\\ 0&1&0&0&1&2\end{array}\right)$ $[x^{6}-(t+13)x^{5}+(9t+63)x^{4}-(27t+145)x^{3}+(31t+165)x^{2}-(11t+87)x+16](x-1)^{t-1}$
$G_{23}$ $\left(\begin{array}[]{cccccc}1&1&0&0&0&0\\ t&t+4&1&1&1&1\\ 0&1&2&1&0&0\\ 0&1&1&3&1&0\\ 0&1&0&1&3&1\\ 0&1&0&0&1&2\end{array}\right)$ $[x^{4}-(t+11)x^{3}+(6t+37)x^{2}-(7t+47)x+20](x-3)(x-1)^{t}$
$G_{26}$ $\left(\begin{array}[]{cccccc}1&1&0&0&0&0\\ t&t+2&1&0&0&1\\ 0&1&2&1&0&0\\ 0&0&1&3&1&1\\ 0&0&0&1&2&1\\ 0&1&0&1&1&3\end{array}\right)$ $[x^{5}-(t+12)x^{4}+(9t+51)x^{3}-(24t+94)x^{2}+(19t+71)x-16](x-1)^{t}$
$G_{27}$ $\left(\begin{array}[]{cccccc}1&1&0&0&0&0\\ t&t+2&0&1&0&1\\ 0&0&1&1&0&0\\ 0&1&1&4&1&1\\ 0&0&0&1&2&1\\ 0&1&0&1&1&3\end{array}\right)$ $[x^{6}-(t+13)x^{5}+(10t+61)x^{4}-(31t+135)x^{3}+(35t+150)x^{2}-(12t+80)x+16](x-1)^{t-1}$
$G_{28}$ $\left(\begin{array}[]{cccccc}1&1&0&0&0&0\\ t&t+3&1&1&0&1\\ 0&1&2&1&0&0\\ 0&1&1&4&1&1\\ 0&0&0&1&2&1\\ 0&1&0&1&1&3\end{array}\right)$ $[x^{6}-(t+15)x^{5}+(11t+84)x^{4}-(40t+228)x^{3}+(58t+319)x^{2}-(29t+221)x+60](x-1)^{t-1}$
$G_{29}$ $\left(\begin{array}[]{ccccc}1&1&0&0&0\\ t&t+3&1&0&2\\ 0&1&2&1&0\\ 0&0&1&3&2\\ 0&1&0&1&4\end{array}\right)$ $[x^{5}-(t+13)x^{4}+(9t+59)x^{3}-(23t+115)x^{2}+(16t+92)x-24](x-2)(x-1)^{t-1}$
$G_{30}$ $\left(\begin{array}[]{ccccc}1&1&0&0&0\\ t&t+3&0&1&2\\ 0&0&1&1&0\\ 0&1&1&4&2\\ 0&1&0&1&4\end{array}\right)$ $[x^{5}-(t+13)x^{4}+(9t+57)x^{3}-(21t+107)x^{2}+(10t+86)x-24](x-2)(x-1)^{t-1}$
$G_{31}$ $\left(\begin{array}[]{ccccc}1&1&0&0&0\\ t&t+4&1&1&2\\ 0&1&2&1&0\\ 0&1&1&4&2\\ 0&1&0&1&4\end{array}\right)$ $[x^{4}-(t+12)x^{3}+(7t+43)x^{2}-(8t+56)x+24](x-3)(x-2)(x-1)^{t-1}$
$G_{32}$ $\left(\begin{array}[]{cccccc}1&1&0&0&0&0\\ t&t+2&1&1&0&0\\ 0&1&3&1&0&1\\ 0&1&1&3&1&0\\ 0&0&0&1&2&1\\ 0&0&1&0&1&2\end{array}\right)$ $[x^{4}-(t+10)x^{3}+(7t+32)x^{2}-(11t+39)x+16](x^{2}-3x+1)(x-1)^{t-1}$
$G_{33}$ $\left(\begin{array}[]{cccc}1&1&0&0\\ t&t+2&0&2\\ 0&0&2&2\\ 0&1&2&3\end{array}\right)$ $[x^{3}-(t+8)x^{2}+(5t+17)x-2t-10](x-3)(x-2)(x-1)^{t-1}x$
$G_{34}$ $\left(\begin{array}[]{ccccc}1&1&0&0&0\\ t&t+3&1&2&0\\ 0&1&3&2&0\\ 0&1&1&3&1\\ 0&0&0&2&2\end{array}\right)$ $[x^{5}-(t+12)x^{4}+(8t+49)x^{3}-(17t+86)x^{2}+(8t+64)x-16](x-3)(x-1)^{t-1}$
$G_{35}$ $\left(\begin{array}[]{ccc}1&1&0\\ t&t+4&4\\ 0&1&5\end{array}\right)$ $[x^{3}-(t+10)x^{2}+(5t+25)x-16](x-3)^{2}(x-1)^{t}$
$G_{36}$ $\left(\begin{array}[]{cccccc}1&1&0&0&0&0\\ t&t+2&1&1&0&0\\ 0&1&4&1&1&1\\ 0&1&1&3&1&0\\ 0&0&1&1&3&1\\ 0&0&1&0&1&2\end{array}\right)$ $[x^{6}-(t+15)x^{5}+(12t+84)x^{4}-(48t+228)x^{3}+(77t+319)x^{2}-(41t+221)x+60](x-1)^{t-1}$
$G_{37}$ $\left(\begin{array}[]{cccc}1&1&0&0\\ t&t+2&0&2\\ 0&0&4&2\\ 0&1&2&3\end{array}\right)$ $[x^{4}-(t+10)x^{3}+(7t+29)x^{2}-(8t+28)x+8](x-3)(x-2)(x-1)^{t-1}$
$G_{38}$ $\left(\begin{array}[]{cccc}1&1&0&0\\ t&t+2&2&0\\ 0&1&5&2\\ 0&0&2&2\end{array}\right)$ $[x^{3}-(t+9)x^{2}+(6t+18)x-8](x-3)(x-2)(x-1)^{t}$
$G_{39}$ $\left(\begin{array}[]{ccccc}1&1&0&0&0\\ t&t+3&1&2&0\\ 0&1&4&2&1\\ 0&1&1&3&1\\ 0&0&1&2&3\end{array}\right)$ $[x^{5}-(t+14)x^{4}+(10t+68)x^{3}-(28t+146)x^{2}+(23t+139)x-48](x-3)(x-1)^{t-1}$
$G_{40}$ $\left(\begin{array}[]{ccccc}1&1&0&0&0\\ t&t+3&2&1&0\\ 0&1&5&1&1\\ 0&1&2&3&0\\ 0&0&2&0&2\end{array}\right)$ $[x^{5}-(t+14)x^{4}+(10t+67)x^{3}-(27t+142)x^{2}+(20t+136)x-48](x-3)(x-1)^{t-1}$
$G_{41}$ $\left(\begin{array}[]{cccc}1&1&0&0\\ t&t+4&2&2\\ 0&1&5&2\\ 0&1&2&3\end{array}\right)$ $[x^{4}-(t+13)x^{3}+(8t+51)x^{2}-(11t+75)x+36](x-3)^{2}(x-1)^{t-1}$
$G_{42}$ $\left(\begin{array}[]{cccc}1&1&0&0\\ t&t+2&2&0\\ 0&1&5&2\\ 0&0&2&4\end{array}\right)$ $[x^{4}-(t+12)x^{3}+(9t+43)x^{2}-(16t+56)x+24](x-3)(x-2)(x-1)^{t-1}$
$G_{43}$ $\left(\begin{array}[]{ccccc}1&1&0&0&0\\ t&t+3&2&1&0\\ 0&1&5&1&1\\ 0&1&2&4&1\\ 0&0&2&1&3\end{array}\right)$ $[x^{4}-(t+13)x^{3}+(9t+51)x^{2}-(15t+75)x+36](x-3)^{2}(x-1)^{t-1}$
$G_{44}$ $\left(\begin{array}[]{ccc}1&1&0\\ t&t+4&4\\ 0&1&7\end{array}\right)$ $[x^{3}-(t+12)x^{2}+(7t+35)x-24](x-3)^{3}(x-1)^{t-1}$